# Extrinsic Calibration of Multiple LiDARs for a Mobile Robot based on Floor Plane And Object Segmentation

### v1

@@ Abstract
Mobile robots equipped with multiple light detection and ranging (LiDARs) and capable of recognizing their surroundings are increasing due to the minitualization and cost reduction of LiDAR. This paper proposes a target-less extrinsic calibration method of multiple LiDARs with non-overlapping field of view (FoV). The proposed method uses accumulated point clouds of floor plane and objects while in motion. It enables accurate calibration with challenging configuration of LiDARs that directed towards the floor plane, caused by biased feature values. Additionally, the method includes a noise removal module that considers the scanning pattern to address bleeding points, which are noises of significant source of error in point cloud alignment using high-density LiDARs. Evaluations through simulation demonstrate that the proposed method achieved higher accuracy extrinsic calibration with two and four LiDARs than conventional methods, regardless type of objects. Furthermore, the experiments using a real mobile robot has shown that our proposed noise removal module can eliminate noise more precisely than conventional methods, and the estimated extrinsic parameters have successfully created consistent 3D maps.
@@ I Introduction
A light detection and ranging (LiDAR) is recognized as a critical sensor in autonomous driving technology due to its high accuracy observation, and it has been advancing in miniaturization and cost reduction. Consequently, the incorporation of LiDARs into mobile robots is increasing. The spreading of high-density LiDARs [1] is also enhancing the perception capabilities of mobile robots. However, it is difficult for a single LiDAR to cover the entire surroundings of a mobile robot, including the travel surface. It is common to mount multiple LiDARs in different directions, such as front and back or left and right, as shown in Fig. 1(a). In the future, it is expected that widespread adoption of these mobile robots equipped with multiple LiDARs.
@@
As a solution to these problems, methods have been proposed that create a common FoV among sensors by accumulating point clouds while in motion [4][5] similar to Fig. 1(b). These methods minimize observation errors between LiDARs by using the entire set of environmental features such as planes and line features accumulated during movement. However, for robots that travel on the ground, LiDARs are often mounted to detect the floor surface, resulting in observations that are predominantly of the floor shown in Fig. 1(c). Consequently, the features obtained may be insufficient, or the floor features may become dominant, leading to unsuccessful extrinsic calibration.
@@
Furthermore, as reported in this research [6], high-density LiDARs are known to exhibit prominent bleeding points, which are generated as it extends from the edge of the object. This noise is caused by the mixing of pulses hitting both foreground and background objects. These noises are a major source of error in extrinsic calibration. Our method includes a module that can remove noise more accurately compared to conventional methods, based on a very simple concept.
@@ II-1 Multiple LiDAR calibration
To solve these problems, methods have been proposed that collect environmental feature while in motion to create a common FoV among LiDARs with no FoV overlap [4]. These methods perform global optimization of a graph composed of extrinsic parameters of each LiDAR and movement trajectories to ensure the consistency of the collected point cloud data. Studies [5][14] have been modified to extract plane and line segment features by adaptive size voxels to improve computational efficiency.
However, these methods needs environments that have sufficient and various direction of features to constrain the 6DoF, such as wall and ceiling. When applying these methods to ground-traveling robots as shown in Fig. 1(a), the point cloud data obtained consists only of the ground and a few objects, leading to insufficient features or dominance of floor point clouds, which can result in unsuccessful extrinsic calibration. To address these issues, our method pre-segment the floor plane and object point clouds, and adopt a two-step estimation process that utilizes each point cloud.
@@ IV-A1 Settings
We created a mobile platform equipped with multiple LiDARs that have non-overlapping FoV in the simulation. The configuration of the multiple LiDARs and simulation environement are shown in Fig. 5. We used the Livox Mid-70 for the LiDARs222https://www.livoxtech.com/mid-70. The LiDARs rotated in place and stoped repeatedly, accumulating data at a total of 16 poses.
@@ IV-B1 Settings
The robot was moved remotely to collect data, and extrinsic calibration was performed for each method. Similar to the simulation, the initial position of the extrinsic parameters was randomly perturbed with a maximum noise of 0.2 m in translation and 5.0 deg in rotation from design value. For the estimation of the motion trajectory, we used an in-house SLAM device mounted on the robot. Figure 9(a) shows the observation scenes. In the scene 1, boxes that were 0.25 m height and these were placed aroud the robot. The various type of objects were placed around the robot in scene 2.
To evaluate the proposed noise removal module, we applied it to the object point clouds and calculated the precision, recall and F-score. For comparison, we compared these values with traditional methods which are SOR[15] and DSOR[17]. We used the same parameters for all methods and used the same parameter for DSOR and proposed method.
@@ V Conclusion
This paper presented a method for extrinsic calibration of multiple LiDARs with non-overlapping FoV, utilizing the accumulated floor information and object point clouds obtained while in motion. Our method pre-segment the floor plane and object point clouds, and adopt a two-step estimation process that utilizes each point cloud. It enables accurate extrinsic calibration for LiDARs that directed towards the ground plane and observation is primarily composed of the floor, a scenario where traditional methods typically struggle. Furthermore, our approach incorporates a noise removal module that considers the scanning patterns to effectively reduce the noise. Simulation-based evaluations have shown that our method can correct randomly given noise by up to 92 % when observing the floor with two or four LiDARs, using various objects. Additionally, the method has been verified for accuracy with actual hardware, demonstrating that our noise removal module outperforms traditional methods in noise reduction, and the estimated extrinsic parameters contribute to the successful creation of consistent 3D maps.
As the spreading of mobile robots continues, it is anticipated that robot inspection facilities will be established, similar to automotive inspection facilities. In such an environment requiring frequent calibration, our proposed method, which does not require a special calibration equipments, is effective in terms of both accuracy and cost. We wish to expand the system to include multimodal devices mounted on robots in our future work.

### v2

@@ Abstract
The utilization of mobile robots equipped with multiple light detection and ranging (LiDAR) sensors, capable of perceiving their surroundings, is on the rise due to the miniaturization and cost reduction of LiDAR technology. This paper introduces a target-less extrinsic calibration method for multiple LiDARs with non-overlapping fields of view (FoV). The proposed method leverages accumulated point clouds of the floor plane and objects obtained during robot motion. It enables accurate calibration, even in challenging configurations where LiDARs are directed towards the floor plane, which can introduce biased feature values. Additionally, the method incorporates a noise removal module that takes into account the scanning pattern to address bleeding points, which are significant sources of error in point cloud alignment when using high-density LiDARs. Evaluations conducted through simulation demonstrate that the proposed method achieves higher accuracy in extrinsic calibration with two and four LiDARs compared to conventional methods, regardless of the type of objects. Furthermore, experiments conducted using a real mobile robot validate the effectiveness of our proposed noise removal module in precisely eliminating noise compared to conventional methods. The estimated extrinsic parameters successfully contribute to the creation of consistent 3D maps.
@@ I Introduction
A light detection and ranging (LiDAR) is recognized as a critical sensor in autonomous driving technology due to its high accuracy observation, and it has been advancing in miniaturization and cost reduction. Consequently, the incorporation of LiDARs into mobile robots is increasing. The spreading of high-density LiDARs [1] is also enhancing the perception capabilities of mobile robots. However, it is difficult for a single LiDAR to cover the entire surroundings of a mobile robot, including the travel surface. It is common to mount multiple LiDARs in different directions, such as front and back or left and right, as shown in Fig. 1(a). In the future, it is expected that widespread adoption of these mobile robots equipped with multiple LiDARs. With the proliferation of such mobile robots, there arises a need for robot inspection sites around the world to calibration at the time of shipment, time of deterioration over time, after overhauls and so on.
@@
To avoid additional tasks such as limitations of the Field of View (FoV) and the preparation of specific components, methods have been proposed that create a common FoV among sensors by accumulating point clouds while in motion [4][5] similar to Fig. 1(b). These methods minimize observation errors between LiDARs by using the entire set of environmental features such as planes and line features accumulated during movement. However, for robots that travel on the ground, LiDARs are often mounted to detect the floor surface, resulting in observations that are predominantly of the floor shown in Fig. 1(c). Consequently, the features obtained may be insufficient, or the floor features may become dominant, leading to unsuccessful extrinsic calibration.
@@
Furthermore, as reported in this research [6], high-density LiDARs are known to exhibit prominent bleeding points, which are generated as they extend from the edge of the objects. This noise is caused by the mixing of pulses hitting both foreground and background objects. These noises are a major source of error in extrinsic calibration. Our method includes a module that can remove noise more accurately compared to conventional methods, based on a very simple concept.
@@ II-1 Multiple LiDAR calibration
To solve these problems, methods have been proposed that collect environmental features while in motion to create a common FoV among LiDARs with no FoV overlap [4]. These methods perform global optimization of a graph composed of extrinsic parameters of each LiDAR and movement trajectories to ensure the consistency of the collected point cloud data. Studies [5][14] have been modified to extract plane and line segment features by adaptive size voxels to improve computational efficiency, and achieves an average translational error of 0.01m and a rotational error of less than 0.1 degrees.
However, these methods needs environments that have sufficient and various direction of features to constrain the 6DoF, such as wall and ceiling. When applying these methods to ground-traveling robots as shown in Fig. 1(a), the point cloud data obtained consists only of the ground and a few objects, leading to insufficient features or dominance of floor point clouds, which can result in unsuccessful extrinsic calibration. To address these issues, our method pre-segment the floor plane and object point clouds, and adopt a two-step estimation process that utilizes each point cloud, and achieves comparable accuracy to above existing methods.
@@ IV-A1 Settings
We created a mobile platform equipped with multiple LiDARs that have non-overlapping FoV in the simulation. The configuration of the multiple LiDARs and simulation environement are shown in Fig. 5. We used the Livox Mid-70 for the LiDARs222https://www.livoxtech.com/mid-70. We tested our proposed method using rotational motion in place and evaluated the error against the ground truth of the extrinsic parameters. The platform performed repeated stopping and rotating actions at poses where both the object and the floor surface could be observed, accumulating data at a total of 16 poses.
@@ IV-B1 Settings
The robot was moved remotely to collect data, and extrinsic calibration was performed for each method. Similar to the simulation, the initial position of the extrinsic parameters was randomly perturbed with a maximum noise of 0.2 m in translation and 5.0 deg in rotation from design value. For the estimation of the motion trajectory, we used an in-house SLAM device mounted on the robot. We conducted optimization using the proposed method by utilizing both the SLAM trajectory and initial extrinsic parameters as input. Figure 9(a) shows the observation scenes. In the scene 1, boxes that were 0.25 m height and these were placed aroud the robot. The various type of objects were placed around the robot in scene 2.
To evaluate the proposed noise removal module, we applied it to the object point clouds and calculated the precision, recall and F-score. To obtain ground truth for the box point cloud, we performed plane removal and manually cropped the point cloud based on the size of the box. For comparison, we compared these values with traditional methods which are SOR[15] and DSOR[17]. We used the same parameters for all methods and used the same parameter for DSOR and proposed method.
@@ V Conclusion
This paper proposes a calibration method for multiple LiDARs with non-overlapping fields of view. By leveraging floor information and point clouds obtained during operation, we achieve precise calibration of LiDARs directed towards the ground. Additionally, we incorporate a noise removal module that takes into account the scanning pattern, effectively reducing noise. Through simulation-based evaluations, we demonstrate the ability of our method to correct up to 92% of randomly given noise when observing the floor and various objects using two or four LiDARs. Furthermore, using real hardware, we confirm the superior noise reduction effects of the noise removal module compared to conventional methods, contributing to the consistent creation of a 3D map. Our proposed method is cost-effective and does not require specialized calibration equipment, making it well-suited for future robot inspection facilities. However, it is important to note that our method assumes operation in flat environments and may experience decreased accuracy in irregular terrains such as outdoor settings. In our future research, we plan to focus on expanding the system to include multimodal devices integrated into robots.